OpenStudy (anonymous):

Find a vector function that represents the curve of intersection of the paraboloid z=7x^2+3y^2 and the cylinder y=2x^2. Use the variable t for the parameter. r(t) =

OpenStudy (anonymous):

r(t)=<t,2t^2,7t^2+12t^4>

OpenStudy (anonymous):

suppose x=t then y=2t^2 and z=7t^2+3(2t^2)^2,z=7t^2+12t^4 so r(t)=(t,2t^2,7t^2+12t^4)